OpenStudy (anonymous):

The quadratic x2 + 20x + 100 has what type of roots? A.Two Irrational B.Two Complex C.Two Rational D.One Rational Double

OpenStudy (anonymous):

I got One Rational Double, but i dont think its right and need help!

OpenStudy (anonymous):

It's right. The discriminant equals zero, thus 1 real solution. x=-10...twice

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

welcome

OpenStudy (unklerhaukus):

When you have a parabola(quadratic) that is of this form\[y=ax^2+bx+c\]its discriminant is \(\Delta=b^2−4ac\) There are three cases \(\Delta<0\) will not touch x axis - there will be no real solutions \(\Delta=0\) will touch x axis at one place only - one solution \(\Delta>0\) will touch x axis exactly twice - two solutions in your case \(a=1,b=20,c=100\) \[\Delta=20^2−4(1)(100)=400-400=0\]